Two girls are 60 miles apart. Leaving at the same time, mary will travel 21 miles/hr and sally will travel 15miles/hr. when will the meet. Please help with the formula.
PL

Answer by glabow(165)

Assume you know where they will meet and call it M. How long does it take Mary to get to M? The answer is 21 mi/hr times x hours. So, we have
21x = M
How long does it take Sally to get to M? The answer is that she must travel the distance that takes her to M starting at 60 mi. And that is 60-M. So, we have
15x = 60-M
Substituting the value of M from the first equation into the second we get
15x = 60 - 21x
36x = 60
x = 5/3
This says that after 5/3 of an hour Mary will have travelled 35 miles (21 x 5/3) and
Sally will have travelled 25 miles (15 x 5/3).
Checking: 35 miles + 25 miles = 60 miles.